What is the rate of return for an investor who pays $1,020 for a three-year bond with a 8% coupon rate and sells the bond one year later for the price carrying 8% YTM?
Answer #1

Price after 1 Year is $ 1000 as Coupon Rate and YTM after 1 year is same ( Both are 8%)

Thus If you Invest $ 1020 today, you will receive 1080 [ 1000 as sale Price and 80 as Coupon amount] after 1 Years

FV = PV [ 1+r]^1

1080 = 1020 [ 1+r]^1

1+r = 1080 / 1020

= 1.0588

r = 1.0588 - 1

= 0.0588 i.e 5.88%